OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y=2/3x y=-2/3x+4

OpenStudy (briellshaw):

so the first step is.....

OpenStudy (anonymous):

@K12girlohva

OpenStudy (alivejeremy):

anybody need help

OpenStudy (er.mohd.amir):

put both y equal to each other

OpenStudy (er.mohd.amir):

-2/3 x +4 =2/3 x

OpenStudy (anonymous):

solution 3xy=2...........i y(3x+4)=-2......ii 3xy+4y=-2.......iii substute eqn (i) into eqn (iii) 2+4y=-2 4y=-4 y=-1 from eqn (i) 3xy=2 but y=-1 3x(-1)=2 -3x=2 x=-2/3 the values of (x,y)=(-2/3,-1)

OpenStudy (anonymous):

hey olivia did u get it

OpenStudy (anonymous):

i cant type a fraction in the answer box :(

OpenStudy (er.mohd.amir):

add eqs then 2y =4 y=2 and 2=2/3 x 6=2x x=3 ans (3,2)

OpenStudy (anonymous):

ight thanks

OpenStudy (anonymous):

i do not think if it is correct @moh since the variables x and y they come together as product...only substitution can save us can u solve it from the starting point